USING PATH INTEGRAL MONTE CARLO SIMULATIONS TO STUDY DELOCALIZATION OF DOPANT MOLECULES IN CLUSTERS OF HELIUM-4

\underline{DAVID T. MOORE}, ROGER E. MILLER, Department of Chemistry, University of North Carolina, Chapel Hill, NC 27599; K. BIRGITTA WHALEY, Department of Chemistry, University of California, Berkeley, CA 94270; YONGKYUNG KWON, Department of Physics, Konkuk University, Seoul 143-701 Korea.

Path Integral Monte Carlo (PIMC) simulations are performed on several dopant molecules in clusters of ^{4}He. We investigate the molecular distributions in the cluster as a function of molecular mass and binding strength to helium, and compare our results with the predictions of simple particle-in-a-spherical-box calculations.